OpenStudy (anonymous):

Write the resultant of the two vectors as an ordered pair. <3,6> and <5,-5>. I already know the answer: <8,1>. Is that right?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@nnesha

OpenStudy (anonymous):

@saseal

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

\[\left[\begin{matrix}3 \\ 6\end{matrix}\right] + \left[\begin{matrix}5 \\ -5\end{matrix}\right] = \left[\begin{matrix}8 \\ 1\end{matrix}\right]\]

OpenStudy (anonymous):